Write Mayer’s formula. The molar heat capacities of a gas at constant pressure
and constant volume are 28.8 J mol−1
K−1
and 20.5 J mol−1
K−1
respectively.
Calculate the gas constant.
1
Expert's answer
2018-04-03T10:24:08-0400
Mayer’s formula can be written as follows:
C_p-C_v=R,
here, C_p is the molar heat capacity of a gas at constant pressure, C_v is the molar heat capacity of a gas at constant volume and R is the gas constant.
Then, we get:
R=C_p-C_v=28.8 J/(mol∙K)-20.5 J/(mol∙K)=8.3 J/(mol∙K).